### v3

@@ Abstract
ZKP systems have surged attention and held a fundamental role in contemporary cryptography. Zk-SNARK protocols dominate the ZKP usage, often implemented through arithmetic circuit programming paradigm. However, underconstrained or overconstrained circuits may lead to bugs. Underconstrained circuits refer to circuits that lack the necessary constraints, resulting in unexpected solutions in the circuit and causing the verifier to accept a bogus witness. Overconstrained circuits refer to circuits that are constrained excessively, resulting in the circuit lacking necessary solutions and causing the verifier to accept no witness, rendering the circuit meaningless. This paper introduces a novel approach for pinpointing two distinct types of bugs in ZKP circuits. The method involves encoding the arithmetic circuit constraints to polynomial equation systems and solving polynomial equation systems over a finite field by algebraic computation. The classification of verification results is refined, greatly enhancing the expressive power of the system. We proposed a tool, AC4, to represent the implementation of this method. Experiments demonstrate that AC4 represents a substantial 29% increase in the checked ratio compared to prior work. Within a solvable range, the checking time of AC4 has also exhibited noticeable improvement, demonstrating a magnitude increase compared to previous efforts.
@@
zero knowledge proofscircuit constraintsunderconstrainedoverconstrainedalgebraic computation checker
@@ 1 Introduction
Developers often opt for the zero-knowledge succinct noninteractive argument of knowledge (zk-SNARK) [9] as the protocol for the efficiency of setting up and proving of building a ZKP system, while also adopting arithmetic circuit [39] as the programming paradigm due to its ease of programming. Circom, a domain-specific language, is based on the aforementioned protocol and programming paradigm. It is capable of converting computational statements into constraints, which can be used to generate a zk system with snark tools such as snarkjs. To illustrate, consider the process of declaring a boolean value in circuit programs. In high-level languages, a boolean value $b$ can be declared directly. However, in circuit languages such as Circom, a constraint must be explicitly defined to limit its possible values, such as the constraint $b\times(b-1)=0$ for $b$. This transformation cannot be automatically performed by existing compilers or libraries. Thus, developers are compelled to express the computations with polynomial equations.
@@
The limitation of polynomial constraint systems gives rise to two problems. The first problem is the underconstrained problem [37], which occurs when multiple outputs can fulfill the same system of equations with identical input values. This situation arises when the system of polynomial equations does not yield a mathematical function on the finite field, rendering the circuit underconstrained. As a result, malicious users could generate fake witnesses, deceive the verifier, and potentially steal the cryptocurrency [35, 14]. The second problem is the overconstrained problem [37], which occurs when no outputs can satisfy the system of polynomial equations or the constraint system is inconsistent for a given input. This problem may give rise to a useless circuit, which in turn opens up the possibility of a denial of service attack. It is obvious that the underconstrained circuits could bring more severe results than the overconstrained circuits.
Identifying underconstrained and overconstrained problems can be achieved by encoding their properties into a logical formula and utilizing a Satisfiability Modulo Theories (SMT) solver to solve them directly. Picus [37] is the first work that makes a valuable effort to employ an SMT solver for investigating the underconstrained cases of ZKP circuits with static program analysis. However, the overconstrained issue still lacks resolution. Nevertheless, the efficiency of SMT solvers brings an obstacle to SMT-based tools when solving large circuits. SMT-based tools struggle to verify constrained circuits within a limited time, even for linear ones.
@@
To this end, an investigation of the constraints on Circom circuit programs is carried out. By reconsidering the formalization paradigm, we model underconstrained and overconstrained problems as algebraic inquiries into whether a quadratic equations system over a finite field has multiple solutions. An algebraic approach is a more intuitive and efficient way to proceed with the realization that the highest degree of the quadratic equations system of equations not exceeding two. Moreover, to support our motivation, we analyzed over 100 popular Circom circuit programs and observed that nearly 70% of them are linear while efficient algorithms exist for solving linear systems of equations. Therefore, after categorizing constraint problems into diverse types according to their maximal degree, we employ a unified arithmetic solution framework to solve them separately.
@@
  • Modeling Circom circuits as polynomial systems over finite fields, we categorize them and propose specialized methods for each category.
  • We develop $\text{AC}^{4}$, a complete and sound tool for verifying circuit programs with the framework above.
  • We evaluate the effectiveness through the solved ratio and solving speed of circuit programs. Our approach solved 65% of previously unsolved examples, a 29% increase over prior work, and significantly improved solving speed.
@@ 2 Background
Zk-SNARK, the dominant ZKP protocol in mainstream tools, frequently uses arithmetic circuits for programming. The verification efforts focus on Circom circuits, which leverage both zk-SNARK and arithmetic circuits.
Zk-SNARKs. Non-interactive ZKPs, such as zk-SNARKs [8], allow verifiers to be assured of a statement’s truth without gaining any knowledge beyond its validity [21]. Blum et al. demonstrated that computational zero-knowledge can be achieved without interaction by sharing a common random string between the prover and the verifier [13, 48]. The non-interactive protocol conveys the proof through a message from the prover to the verifier, ensuring efficiency. zk-SNARKs reign supreme in zk systems due to lightning-fast verification of even large programs, using tiny proofs (mere hundreds of bytes) within milliseconds.
@@
Circom. Circom is a compiler written in Rust [33] for compiling circuits written in the Circom language, a SNARK-based language that uses arithmetic circuits as its programming paradigm. The compiler outputs circuit constraints and all the necessary information for computing different ZK proofs. In Circom, programmers can define constraints using the operators ======= = = for equality, <<--< - - or >-->- - > for signal assignment, and <==absent<==< = = or ==>==>= = > for simultaneous assignment. Constraints in Circom must be in the form of $A\times B+C=0$, where $A$, $B$, and $C$ are linear combinations of signals. After compiling circuit to r1cs format and computing the witness with the input, users can utilize the snarkjs tool to generate and validate a proof for the specified input. The verifier will receive the proof if the witness is valid, otherwise, the verifier will reject it.
We address Circom circuit constraint verification by extending arithmetic circuit definitions with constraints.
Definition 2 (Constraint)
@@
An arithmetic circuit with constraint with constraints can be defined as a tuple $\langle C(\mathcal{X}),E\rangle$, where $C(\mathcal{X})$ is an arithmetic circuit. $\mathcal{X}$ is split into two disjoint sets $K$ and $U$. $K$ represents known input variables, and $U=T\cup O$ represents unknown variables including the non-constant variables including temporal variables $T$ and output variables $O$. $E$ includes all the constraints of the circuit $C(\mathcal{X})$.
@@ 3 Problem Statements
Detecting overconstrained or underconstrained circuits in Circom boils down to an algebraic inquiry into the number of solutions for a quadratic equations system over a finite field. This inherent connection naturally leads us to model Circom circuits as such systems and leverage algebraic methods for analysis. Moreover, the design principle of R1CS forces Circom constraints under quadratic computation, resulting a maximum degree of 2 for the equivalent equations. Our method follows the classification-and-discussion scheme. Our strategy categorizes circuits by degree and applies specialized solutions within each category, leading to significant efficiency and scalability gains.
@@ 3.1 Constraint Problems
The construction of arithmetic circuits is plagued by two critical challenges: overconstrained and underconstrained circuits. Programmer-authored circuits, even from seasoned ZK developers, often bear inadvertently miscalibrated constraints, leading to these issues. While completely eradicating these issues remains a persistent hurdle, current approaches fall short of effectively addressing them. This subsection tackles this double-edged sword head-on, proposing a novel approach to ensure the construction of robust and reliable arithmetic circuits.
Overconstrained bugs. As described in the work by Labs et al. [27], overconstrained circuits are unable to uphold the completeness of zero-knowledge proof. A Circom program is deemed overconstrained when its corresponding polynomial constraints are overconstrained, indicating that the circuit cannot produce any valid outputs regardless of the inputs provided. This inconsistency in the constraint system for a given input characterizes the state of being overconstrained. Despite their deceptive simplicity in concept and the misconception that they are less critical than other types of attacks, denial of service (DoS) attacks are often underappreciated regarding their security implications. Notably, DoS attacks require neither high skill nor deep knowledge,, contributing to their low cost for malicious attackers as highlighted by Zcash [30].
Underconstrained bugs. Recent studies have discovered that underconstrained bugs in arithmetic circuits can lead to various security threats, including the ability for attackers to forge signatures, steal user concurrency, or create counterfeit crypto-coins. A Circom program is classified as underconstrained when the corresponding polynomial constraints are also underconstrained. This condition arises when there exists an input $in$ for which multiple distinct outputs $out_{1},out_{2},\ldots$ can be produced, and the circuits $C(in,out_{1}),C(in,out_{2}),\ldots$ are all considered valid. In essence, an underconstrained circuit program may unintentionally permit the acceptance of unwanted signals, making it susceptible to exploitation by malicious entities.
@@ 3.2 Categories of Elements
In our method, categorizations are necessary to account for the various types of variables present in arithmetic circuits. Some variables are determined by external inputs, while others depend on other variables, necessitating the need to classify them accordingly. Aligning with Circom’s degree-2 limit, it is pertinent to categorize the variables based on the degree. Furthermore, due to the computational complexity of verifying certain circuits, we are limited to providing coarse-grained results, thereby highlighting the importance of categorizing the outcomes as well.
@@
Category of Variables. The variables in a Circom circuit consist of the input signals $I$, temporary signals $T$, and output signals $O$. The user provides the values for $I$, which can be regarded as known values $K$. On the other hand, the variables in $O\cup T$, whose values are not constant, are regarded as unknown variables $U$ within the constraints. The key issue of the constraint problem is identifying the unique characteristics of the $O$ variables.
@@
Category of Equations and Circuits. The constraint $e$ is an algebraic polynomial equation whose degree is under 2 over a finite field $F$, while the circuits are a system of algebraic polynomial equations whose degree is under 2 over a finite field, with their generator variables being $U$.
@@
The three types of equations (constraints) to consider are precise linear equations, $K$-coefficient linear equations, and quadratic equations. A precise linear equation is defined by having linear generator variables with constant coefficients. Conversely, a $K$-coefficient linear equation includes at least one generator variable with a coefficient represented by the known variables $K$, and none of the generator variables’ coefficients are represented by the unknown variables $U$. Lastly, a quadratic equation is distinguished by containing at least one generator variable with a coefficient represented by the unknown variables $U$.
@@
The system of equations, namely circuits, can be categorized into three types based on the constraints outlined above: precise linear circuits, $K$-coefficient circuits, and quadratic circuits. A precise linear circuit consists solely of precise linear equations, while a $K$-coefficient circuit must contain at least one $K$-coefficient linear equation and no quadratic equations. Finally, a quadratic circuit should include at least one quadratic equation.
Category of Results. As shown in Fig. 1, $\text{AC}^{4}$ yields six result types: algebraic overconstrained, precisely overconstrained, algebraic exactconstrained, precisely exactconstrained, precisely underconstrained, and unknown. The algebraic overconstrained and algebraic exactconstrained results are collectively referred to as algebraically solved results, while precisely overconstrained and precisely exactconstrained results are collectively referred to as precisely solved results.
The verification problem is refined to create a more detailed classification system for the output, based on constraints safety metrics. As for the unknown cases that are not fully solved by the solver, they are categorized into two intermediate cases separated by constraint stress. The two intermediate cases, algebraically exact constrained and algebraically over constrained, provided more detailed information for analysis. Besides this category, this method unlocks partial information about unsolvable circuits and speeds things up.
Fig. 1 also provides a cross-reference between our result category and the typical ground truth of constraints. Among those 5 practical output types except unknown  3 precise categories of them: precisely underconstrained, precisely exactconstrained, and precisely overconstrained, strictly maps into the underconstrained, exactconstrained, overconstrained respectively. The 2 medium status algebraic exactconstrained and algebraic overconstrained, caused by potential reduced-rank phenomenon while simplifying the equation matrix in some cases. Our approach cannot ascertain equation system solution counts, reflected in matrix rank. Cases labeled with algebraic exactconstrained may lie in underconstrained or exactconstrained of ground truth, while cases labeled with algebraic overconstrained occur in exactconstrained or overconstrained.
Algebraically overconstrained indicates no algebraic solutions, but special inputs can shift results to exactconstrained or underconstrained. The circuit is algebraically exactconstrained, indicating it has exactly one solution algebraically. However, Special inputs can trigger underconstrained states.
Unknown results may stem from circuit size, computational complexity, or the inherent NP-hardness of solving quadratic equations over finite fields.
By combining the algebraic exactconstrained and precisely exactconstrained into a synthesized state, namely general exactconstrained, cases in this state are never overconstrained, though they could fall in underconstrained. Similarly, the general overconstrained state provides a guarantee for cases to avoid underconstrained state. We introduce a relaxed constraint property metric that efficiently filters out under/overconstrained cases.
Figure 1: The category of results from AC4.
@@ 3.3 Problem Reduction
For each circuit $A=\langle C(X),e\rangle\in\langle C(\mathcal{X}),E\rangle$, there exists $A^{\prime}\in\mathbb{F}[X_{1},\ldots,X_{k}]$, such that $A^{\prime}=\mathbb{M}(A)$ , where $\mathbb{M}$ is a bijection function from a circuit to quadratic polynomials over finite fields.

### v4

@@ Abstract
Zero-knowledge proof (ZKP) systems have surged attention and held a fundamental role in contemporary cryptography. Zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) protocols dominate the ZKP usage, implemented through arithmetic circuit programming paradigm. However, underconstrained or overconstrained circuits may lead to bugs. The former refers to circuits that lack the necessary constraints, resulting in unexpected solutions and causing the verifier to accept a bogus witness, and the latter refers to circuits that are constrained excessively, resulting in lacking necessary solutions and causing the verifier to accept no witness. This paper introduces a novel approach for pinpointing two distinct types of bugs in ZKP circuits. The method involves encoding the arithmetic circuit constraints to polynomial equation systems and solving them over finite fields by the computer algebra system. The classification of verification results is refined, greatly enhancing the expressive power of the system. A tool, AC4, is proposed to represent the implementation of the method. Experiments show that AC4 demonstrates a increase in the checked ratio, showing a 29% improvement over Picus, a checker for Circom circuits, and a 10% improvement over halo2-analyzer, a checker for halo2 circuits. Within a solvable range, the checking time has also exhibited noticeable improvement, demonstrating a magnitude increase compared to previous efforts.
@@
zero-knowledge proofscircuit constraintsunderconstrainedoverconstrainedalgebraic computation checker
@@ 1 Introduction
Developers frequently choose the zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) [12] protocol for its efficiency in setting up and proving within a zero-knowledge proof (ZKP) system. Additionally, they adopt arithmetic circuits [48] as the programming paradigm due to their simplicity in representing computational problems. Both Circom and halo2 are built upon this protocol and programming paradigm. To illustrate this, consider the process of declaring a boolean value in circuit programs. In high-level languages, a boolean value $b$ can be declared directly. However, in circuit languages, a constraint must be explicitly defined to limit its possible values. For instance, the constraint $b\times(b-1)=0$ ensures that $b$ can only be 0 or 1. This transformation cannot be automatically performed by existing comilers or libraries, necessitating that developers express computations using polynomial equations.
@@
The limitation of polynomial constraint systems gives rise to two problems. The first problem is the underconstrained problem [46], which occurs when multiple outputs can fulfill the same system of equations with identical input values. This situation arises when the system of polynomial equations does not yield a mathematical function on the finite field, rendering the circuit underconstrained. As a result, malicious users could generate fake witnesses, deceive the verifier, and potentially steal the cryptocurrency [4, 19]. The second one is the overconstrained problem [46], which occurs when no outputs can satisfy the system of polynomial equations or the constraint system is inconsistent for a given input. This issue can result in a non-functional circuit. Consequently, this may lead to overconstrained bugs that cause ZK-verifier servers to reject numerous legitimate requests. Such rejections could potentially trigger Distributed Denial of Service (DDoS) problems. Furthermore, it is evident that underconstrained circuits can produce more severe consequences than their overconstrained counterparts.
@@
To this end, we conducted an investigation into the constraints on Circom or halo2 circuit programs. By reconsidering the formalization paradigm, we modeled underconstrained and overconstrained problems as algebraic inquiries. Specifically, we examined whether a system of polynomial equations over a finite field has multiple solutions. To support our motivation, we analyzed the official library of circom circuit templates provided by iden3 and collected halo2 circuit templates from . Our analysis revealed that nearly 70% of these templates are linear, and efficient algorithms exist for solving linear systems of equations. Consequently, we categorized constraint problems into diverse types based on their maximal degree. We then employed a unified arithmetic solution framework to address each type separately.
@@
  • We proposed a methodology that utilizes computer algebra systems to model circuit programs as polynomial systems over finite fields, with the aim of verifying their correctness.
  • We developed AC4, a highly efficient and practical tool that leverages the aforementioned framework to effectively handle larger circuit programs.
  • We constructed datasets specifically tailored for Circom and halo2 circuits. By comparing the checked ratio and checking speed of AC4 against other tools, we observed substantial improvements in both metrics.
@@ 2 Background
Zk-SNARKs. Non-interactive ZKP protocols, such as zk-SNARKs [11], allow verifiers to be assured of a statement’s truth without gaining any knowledge beyond its validity [27]. Blum et al. demonstrated that computational zero-knowledge can be achieved without interaction by sharing a common random string between the prover and the verifier [15, 58]. The non-interactive protocol conveys the proof through a message from the prover to the verifier, ensuring efficiency. zk-SNARKs reign supreme in ZKP systems due to lightning-fast verification of even large programs, using tiny proofs (mere hundreds of bytes) within milliseconds.
@@
Circuit Languages. The growing field of ZKPs has led to the development of specialized circuit languages designed to create and manage the underlying circuits needed for these proofs. Two notable examples are the Circom [9] and halo2 [59] frameworks.
Circom is a domain-specific language that leverages SNARK technology and employs arithmetic circuits as its programming paradigm. Programmers define circuits using specific operators: ======= = = for equality checks, <<--< - - or >-->- - > for signal assignment, and <==absent<==< = = or ==>==>= = > for simultaneous assignment. These circuits conform to the rank-1 constraint system (R1CS) format. The R1CS format is expressed as $A\times B+C=0$, where $A$, $B$, and $C$ are linear combinations of signals. After compiling the circuit into R1CS format and computing a witness using the input, the snarkjs tool can be employed to generate and verify a ZKP for the specified input.
Halo2 provides a more flexible and scalable method for constructing ZKPs. It employs PLONKish arithmetization [26], which supports recursive proof composition, thereby enhancing scalability. In halo2, circuits are built using polynomials and lookups. This construction allows for more efficient proof generation and verification.
It is reasonable to conclude that arithmetic circuits in both circom and halo2 can be treated as sets of polynomial equations, despite differing in their representations, methods of construction, and evaluation. In particular, halo2 supports constraints that are typically quadratic or higher. The specific implementation and optimization goals determine the exact nature of these constraints. We address circuit constraint verification by extending arithmetic circuit definitions with constraints.
Definition 2 (Constraint in Circom)
@@
A halo2 circuit constraint is more general.
Definition 3 (Constraint in halo2)
A constraint in halo2 is a polynomial equation over a finite field $F_{p}$:
$\sum_{i=0}^{n}\sum_{j=0}^{n}a_{i,j}\cdot x_{i}\cdot x_{j}+\sum_{i=0}^{n}b_{i}\cdot x_{i}+c=0$
where $n$ is the degree of the constraint, $a_{i,j},b_{i},c\in F_{p}$ are the coefficient scalars, and $x_{i}\in F_{p}$ are the variables.
@@
An arithmetic circuit with constraint with constraints can be defined as a tuple $\langle C(\mathcal{X}),E\rangle$, where $C(\mathcal{X})$ is an arithmetic circuit on variable set $\mathcal{X}$. $\mathcal{X}$ is the union of two disjoint sets $K$ and $U$, where $K$ represents known input variables, and $U=T\cup O$ represents unknown variables including the non-constant variables including temporal variables $T$ and output variables $O$. $E$ includes all the constraints of $C(\mathcal{X})$.
Computer algebra. Computer algebra [16] revolves around the manipulation of algebraic expressions and various mathematical entities. This field has spurred the creation of computer algebra systems capable of autonomously tackling a broad spectrum of theoretical and practical mathematical challenges. A contemporary computer algebra system boasts a diverse array of functionalities, encompassing Gröbner bases, cylindrical algebraic decomposition, lattice basis reduction, linear system solving, arbitrary and high-precision arithmetic, interval arithmetic, linear and nonlinear optimization, Fourier transforms, Diophantine equation solving, among others. These capabilities are indispensable for both theoretical investigations and practical applications.
Gröbner basis. A Gröbner basis is a set of multivariate polynomials that has desirable algorithmic properties. Every set of polynomials can be transformed into a Gröbner basis. This process generalizes three familiar techniques: Gaussian elimination for solving linear systems of equations, the Euclidean algorithm for computing the greatest common divisor of two univariate polynomials, and the Simplex algorithm for linear programming. Algorithms on Gröbner basis are widely used in mathematical and verification tools. The computation of a Gröbner basis can be performed using Buchberger’s algorithm [54], while Faugère F5 algorithm [23] is a significant improvement over the former. It introduces the concept of signatures to predict and eliminate unnecessary computations, which can potentially increase the efficiency of Gröbner basis computation.
@@ 3 Problem Statements
Detecting overconstrained or underconstrained circuits in circuit languages boils down to an algebraic inquiry into the number of solutions for a polynomial equations system over a finite field. This inherent connection naturally leads us to model circuits as such systems and leverage algebraic methods for analysis. Our method follows the classification-and-discussion scheme. Our strategy categorizes circuits by degree and applies specialized solutions within each category, leading to significant efficiency and scalability gains.
@@ 3.1 Constraint Problems
The construction of arithmetic circuits is plagued by two critical challenges: overconstrained and underconstrained circuits. Programmer-authored circuits, even from seasoned ZKP developers, often bear inadvertently miscalibrated constraints, leading to these issues. While completely eradicating these issues remains a persistent hurdle, current approaches fall short of effectiveness. This subsection tackles this double-edged sword head-on, proposing a novel approach to ensure the construction of robust and reliable arithmetic circuits.
Overconstrained bugs. As described in the work by Labs et al. [35], overconstrained circuits are unable to uphold the completeness of ZKP. A circuit program is deemed overconstrained when its corresponding polynomial constraints are overconstrained, indicating that the circuit cannot produce any valid outputs regardless of the inputs provided. This inconsistency in the constraint system for a given input characterizes the state of being overconstrained. Despite their deceptive simplicity in concept and the misconception that they are less critical than other types of attacks, denial of service (DoS) attacks are often underappreciated regarding their security implications. Notably, DoS attacks require neither high skill nor deep knowledge,, contributing to their low cost for malicious attackers as highlighted by Zcash [38].
Underconstrained bugs. Recent studies have discovered that underconstrained bugs in arithmetic circuits can lead to various security threats, including the ability for attackers to forge signatures, steal user concurrency, or create counterfeit crypto-coins. A circuit program is classified as underconstrained when the corresponding polynomial constraints are also underconstrained. This condition arises when there exists an input $in$ for which multiple distinct outputs $out_{1},out_{2},\ldots$ can be produced, and the circuits $C(in,out_{1}),C(in,out_{2}),\ldots$ are all considered valid. In essence, an underconstrained circuit program may unintentionally permit the acceptance of unwanted signals, making it susceptible to exploitation by malicious entities.
@@ 3.2 Categories of Elements
In our method, categorization is essential to account for the different types of variables in arithmetic circuits. Variables can be classified based on their source: some are determined by external inputs, while others depend on internal variables. It is also important to categorize variables according to their degree. Due to the computational complexity of verifying certain circuits, we can only provide coarse-grained results. This limitation underscores the importance of categorizing the outcomes as well.
@@
Category of Variables. The variables in a circuit program consist of the input signals $I$, temporary signals $T$, and output signals $O$. The user provides the values for $I$, which can be regarded as known values $K$. On the other hand, the variables in $O\cup T$, whose values are not constant, are regarded as unknown variables $U$ within the constraints. The key issue of the constraint problem is identifying the unique characteristics of the $O$ variables.
@@
Category of Constraints and Circuits. The constraint $e$ is an algebraic polynomial equation whose degree is under 2 over a finite field $F$, while the circuits are a system of algebraic polynomial equations whose degree is under 2 over a finite field, with their generator variables being $U$.
@@
The three types of constraints to consider are precise linear equations, $K$-coefficient linear constraints, and higher constraints. A precise linear constraint is defined by having linear generator variables with constant coefficients. Conversely, a $K$-coefficient linear equation includes at least one generator variable with a coefficient represented by the known variables $K$, and none of the generator variables’ coefficients are represented by the unknown variables $U$. Lastly, a higher equation is distinguished by containing at least one generator variable with a coefficient represented by the unknown variables $U$.
@@
The system of constraints, namely circuits, can be categorized into three types based on the constraints outlined above: precise linear circuits, $K$-coefficient circuits, and higher-order circuits. A precise linear circuit consists solely of precise linear constraints, while a $K$-coefficient circuit must contain at least one $K$-coefficient linear constraint and no higher-order constraints. Finally, a higher-order circuit should include at least one higher-order constraint.
@@ 3.3 Problem Reduction
For each circuit $A=\langle C(X),e\rangle\in\langle C(\mathcal{X}),E\rangle$, there exists $A^{\prime}\in\mathbb{F}[X_{1},\ldots,X_{k}]$, such that $A^{\prime}=\mathbb{M}(A)$ , where $\mathbb{M}$ is a bijection function from a circuit to polynomial equations over finite fields.